40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5178     Entry time: Wed Aug 10 19:18:26 2011
Author: Nicole 
Type: Summary 
Category: SUS 
Subject: Fixed Reflective Photosensors; Recalibrated Photosensor 2 

Thanks to Koji's help, the second photosensor, which was not working, has been fixed. I have re-calibrated the photosensor after fixing a problem with the circuit.  I have determined the new linear region to lie between 7.6 mm and 19.8mm. The slope defining the linear region is -0.26 V/mm (no longer the same as the first photosensor, which is -0.32 V/mm).

 

Here is the calibration plot.

ps2.jpg

ELOG V3.1.3-